1. IN 1990, THERE WERE A NUMBER OF HOUSES IN THE STREET WHRERE BRAD LIVED. TWO YEARS AFTER,5 HOUSES WRE ADDED. SIX DEMOLISHED A YEAR AFTER.IF AT PRESENT THERE WERE 30 HOUSES IN STREET, HOW MANY HOUSES WERE THERE IN 1990
2.A NUMBER WHEN DECREASED BY THE SUM OF 7 AND 8 WILL GIVE YOU 20.WHAT IS THE NUMBER

Answer by rfer(16322)

A)
Run the numbers in reverse order.
30+6=36-5=31 to start
B)
7+8=15+20=35